OpenStudy (anonymous):

@Loveiskey18 After 3 years, Remington earned $390 in simple interest from a CD into which he initially deposited $4000. What was the annual interest rate of the CD? A. 13% B. 6.5% C. 3.25% D. 9.75%

OpenStudy (anonymous):

All you do is subtract 4000-by your options until you get 390. *hint begin with 9.75%*. (4000-9.75% =?)

OpenStudy (trantom):

390/3=130 130/4000=0.0325 3.25%

OpenStudy (trantom):

C

OpenStudy (trantom):

loveiskey18 you CAN do that but it's better and faster to use algebra then guess and check.

OpenStudy (anonymous):

Your answer is incorrect. Do not accuse me of guessing. :)